OpenStudy (firejay5):

Show work!!!!!!!!! 5. A box of that has a mass of 34kg is pushed horizontally across the garage floor with a force of 185N. What is the normal force acting on the box?

OpenStudy (firejay5):

@midhun.madhu1987 I am sure how to solve this one Fn = m * g Fn = (24) (9.8) Fn = 333N Is the above answer all I have to do? What about the 185N?

OpenStudy (firejay5):

***I am not sure

OpenStudy (firejay5):

@midhun.madhu1987 Why do you say 0?

OpenStudy (midhun.madhu1987):

My mistake... i thought of work... it might be 34 * 9.8

OpenStudy (firejay5):

that was my mistake sorry; what about after I multiply 34 * 9.8; what next?

OpenStudy (midhun.madhu1987):

 that is it... i guess